OpenStudy (anonymous):

Wrong answer??? Find the equation of the sphere if one of its diameters has endpoints (-5, -3, -7) and (-4, -1, -4) I got (x+4.5)^2+(y+2)^2+(z+1.5)^2-31.5

OpenStudy (radar):

What did you get for the location of the center of the sphere?

OpenStudy (radar):

What did you get for the radius?

OpenStudy (radar):

\[x _{0}=1/2\]\[y _{0}=-1\]\[z _{0}=-3/2\] Is that the coordinates of the center?

OpenStudy (radar):

\[(x-1/2)^{2}+(y+1)^{2}+(z+3/2)^{2}=258/4\]Definitely different from your answer.